### v1

# Bridging Deep Reinforcement Learning and Motion Planning for Model-Free

Abstract
Deep Reinforcement Learning (DRL) has emerged as a powerful model-free paradigm for learning optimal policies. However, in real-world navigation tasks, DRL methods often suffer from insufficient exploration, particularly in cluttered environments with sparse rewards or complex dynamics under system disturbances. To address this challenge, we bridge general graph-based motion planning with DRL, enabling agents to explore cluttered spaces more effectively and achieve desired navigation performance. Specifically, we design a dense reward function grounded in a graph structure that spans the entire state space. This graph provides rich guidance, steering the agent toward optimal strategies. We validate our approach in challenging environments, demonstrating substantial improvements in exploration efficiency and task success rates. The project website is available at: https://plen1lune.github.io/overcome_exploration/
Index Terms:
Deep Reinforcement Learning, Motion Planning, Sampling-based Method
I Introduction
Recently, the demand for learning control policies increasingly prevalent across a wide range of real-world tasks. Examples include agile drone flight in challenging scenes [1], mobile robot navigation [2, 3, 4], and object manipulation [5]. These tasks often involve high-dimensional sensory inputs and uncertainty in dynamics, where classical control methods are insufficient. For such tasks, DRL has become a dominant framework for obtaining control policies, thanks to its model-free nature that allows agents to learn directly from interactions without relying on explicit dynamics models [6, 7]. This flexibility makes DRL particularly appealing in domains such as robotic manipulation [8, 9], autonomous driving [10], and locomotion [11], where constructing accurate models is often infeasible. In these applications, DRL offers the potential to generalize across tasks and adapt to uncertainties, establishing itself as a key approach for intelligent decision-making under complex conditions.
However, the practical success of DRL remains severely limited by its sample inefficiency, primarily due to inadequate exploration in sparse-reward or high-dimensional settings [12, 13]. Without informative feedback, agents may spend substantial time interacting with irrelevant states. To address this issue, a range of exploration strategies have been proposed. Intrinsic motivation methods reward agents for visiting novel or uncertain states [14, 15], while count-based approaches approximate visitation frequencies using density models [16, 13]. Noise-driven methods, such as entropy regularization [17] or stochastic policy perturbation [18], also remain popular due to simplicity. Some studies decompose the exploration process either by partitioning the task into subtasks [19] or employ temporal logic (TL) to guide exploration [20].While these methods partially enhance sample efficiency, they fall short of resolving the fundamental difficulty of agent exploration in cluttered environments.
Classical motion planning algorithms, such as RRT, provide structured strategies for navigating cluttered environments by exploiting geometric information and long-horizon planning capabilities [21, 22]. However, these approaches are typically integrated with model-based path-tracking control approaches such as Model Predictive Control (MPC) [23]. Model-free path-tracking, which is the focus of this paper, is still an open problem. To address this limitation, Cai et al. [24] proposed a motion planning-guided DRL framework, where waypoints are generated via Rapidly-exploring Random Trees (RRT) to guide agent exploration. This integration provides structured, long-horizon guidance, boosting sample efficiency and enabling faster convergence, especially in maze-like or constrained environments [25]. The agent leverages both classical planning coverage and model-free policy adaptation to explore efficiently. However, the planned trajectories may themselves be infeasible due to environmental constraints, or become invalid under uncertainty [3, 26, 27]. This static nature of the path limits adaptability and robustness, and calls for learning frameworks that can handle dynamic replanning and online corrections. However, in certain real-world tasks, real-time replanning may be impractical due to computational constraints, latency requirements, or limited sensing capabilities, making it essential to develop exploration strategies that are both efficient and lightweight [28, 29, 30].
Contributions: In this work, we propose a novel graph-based framework that enhances RL exploration in cluttered environments and can be seamlessly integrated with a wide range of model-free RL algorithms. Compared to previous work [24], our method achieves more comprehensive exploration of the environment’s state space, allowing agents to better handle challenging scenarios and potential disturbances, while fully harnessing the capabilities of model-free reinforcement learning to learn optimal policies under unknown dynamics. We further provide theoretical guarantee that our exploration strategy preserves the original RL objective and accelerates training convergence. In addition, our framework enables agents to generalize across arbitrary initial states without retraining or policy modification.
Organization: Sec. II mentions the foundations of our approach. In Sec. III, we define the problem and emphasize the challenges. Sec. IV presents our approach for addressing simple goal-reaching tasks via reward design using sampling-based methods over the workspace. Sec. V demonstrate how we generalize the initialization positions. The performance of the proposed method is shown in Sec. VI.
II Related Works
Motion Planning: Sampling-based motion planning methods, such as Rapidly-exploring Random Tree (RRT) [31], RRT* [22], Probabilistic Road Map (PRM) [32], and Rapidly-exploring Random Graph (RRG) [22], are widely used for finding collision-free paths in continuous geometric spaces. Among these, PRM constructs a roadmap by sampling random points in the configuration space, connecting them through feasible edges based on a local planner, and using this graph representation to find paths between specified start and goal points. This method is particularly efficient for multi-query settings, as the precomputed roadmap can be reused across different planning tasks. RRG extends the idea of RRT by maintaining a graph structure that connects all sampled points, thereby ensuring asymptotic optimality while allowing more flexible pathfinding. To take advantage of these graph structures for navigation, shortest-path algorithms such as A* [33] are often employed to compute optimal paths on the graph. A* performs heuristic-guided search to efficiently identify paths that minimize cost, making it a powerful tool for motion planning tasks involving graph-based representations. While these methods provide robust solutions for high-dimensional motion planning, they typically rely on model-based path-tracking controllers to execute the planned trajectories.
Deep Reinforcement Learning: Deep Reinforcement Learning (DRL) has emerged as a powerful model-free paradigm for learning optimal policies in complex decision-making problems [7]. In continuous control tasks, DRL algorithms such as Proximal Policy Optimization (PPO) [18] and Soft Actor-Critic (SAC) [17] have been widely adopted due to their empirical stability and sample efficiency. These algorithms are typically built upon the actor-critic framework, which has become the standard paradigm in modern policy-based reinforcement learning [7, 17]. In this framework, the actor is a parameterized policy that maps states to actions, while the critic commonly estimates the expected return—either in the form of a state-value function or an action-value function—to guide the improvement of the actor. This separation enables the actor to focus on policy optimization while leveraging the critic to provide accurate feedback based on the agent’s experience, thereby facilitating more effective and stable learning dynamics.
III Problem Formulation
System Disturbance: We consider a continuous-time dynamical system $S$ with general unknown dynamics, whose evolution is described by:
$\frac{ds}{dt}=\dot{s}=f(s,a,d)$ (1)
where $s\in\mathcal{S}\subseteq\mathbb{R}^{n}$ represents the state variables, $a\in\mathcal{A}\subseteq\mathbb{R}^{m}$ denotes the control inputs, and $d$ is the disturbance function.
MDPs: Markov Decision Processes (MDPs) provide a mathematical framework for modeling sequential decision-making problems in stochastic environments. In the discounted setting, we consider a tuple $(\mathcal{S},\mathcal{A},p,r,\rho_{0},\gamma)$, where $\mathcal{S}$ denotes a continuous state space, $\mathcal{A}$ denotes an continuous action space, $p:\mathcal{S}\times\mathcal{S}\times\mathcal{A}\to[0,1]$ denotes a transition function, $r:\mathcal{S}\times\mathcal{A}\to[r_{\min},r_{\max}]$ be a bounded reward function. $\rho_{0}:\mathcal{S}\to[0,1]$ denotes an initial distribution, and $\gamma\in(0,1]$ is a discount factor. At the beginning of the process, the agent’s initial state $s_{0}$ is sampled from the initial state distribution $\rho_{0}$. At each time step $t$, the agent observes the current state $s_{t}$, and selects an action $a_{t}$ based on its policy $\pi:\mathcal{S}\to\mathcal{P}(\mathcal{A})$, which is a mapping from states to probability distributions on $\mathcal{A}$. The agent samples an action according to the probability distribution, i.e., $a_{t}\sim\pi(\cdot|s_{t})$. Once the agent takes the action $a_{t}$, the environment responds by transitioning the agent to a new state $s_{t+1}$, governed by the transition function $p$, i.e., $s_{t+1}\sim p(\cdot|s_{t},a_{t})$. After that, it receives a scalar reward $r_{t+1}$ from the environment. This reward reflects the immediate benefit or cost of taking action $a_{t}$ in state $s_{t}$. To balance immediate and future rewards, the agent uses the discount factor $\gamma$, which ensures that rewards received earlier in time are more valuable than those in the distant future. The total discount reward, also called the return, is calculated as $G_{t}=\sum_{k=0}^{\infty}\gamma^{k}r(s_{t+k},a_{t+k})$, where $G_{t}$ denotes the return at the time step $t$. Given the start state $s$ and policy $\pi$, we define the state value function as $V_{\pi}(s)=\mathbb{E}_{\pi,p}\left[G_{t}|s_{t}=s\right]$.
Figure 1: Comparison between previous method and graph-based method. We implemented both RRT (left) and RRG (right) to solve the same goal-reach task. As shown, the gray areas represent known obstacles, and the red cross indicate potential disturbances that may lead to failure transitions.
Problem.
Consider a navigation task in a cluttered environment with unknown system dynamics. The goal is to learn the optimal controllers $\pi^{*}$ that enable safe navigation of goal-reaching tasks.
This problem concerns finding a feasible solution in challenging environments. While standard MDP theory guarantees the existence of an optimal policy under mild conditions, learning such a policy via RL remains difficult in sparse-reward or long-horizon settings. These challenges are further exacerbated in practice by the clutter of the environment. Our work specifically addresses these issues by enhancing exploration through a graph-based guidance framework.
Example 1.
We compared our approach with methods that use motion planning-generated paths to guide the training of reinforcement learning agents [24] (see Figure 1). To illustrate the limitations, we deliberately introduced an infeasibility (indicated by a red cross) along the planned geometric path to the goal. The path-based method (left) fails to manage such disturbances, as it relies on a single, fixed trajectory without considering alternative routes, resulting in failure. Consequently, using waypoints from this path as guidance also fails. In contrast, our approach employs a graph that covers the entire workspace (right), providing more robust guidance.
IV Overcoming Exploration
In Section IV, we introduce the graph-based reward and our design enabling arbitrary starting points.
IV-A Geometric Graph-based Motion planning Methods
While our implementation employs RRG for demonstration purposes, the proposed framework is fundamentally agnostic to the choice of graph construction methods. In particular, it accommodates both sampling-based planners, such as RRG [22] and PRM [32], and heuristic-guided search algorithms, notably A* [33]. The RRG incrementally constructs an undirected graph $G=(V,E)$ in the configuration space $\mathcal{X}$, where $V$ is the set of vertices representing sampled points, and $E$ consists of edges that connect pairs of vertices via collision-free paths. Compared to tree-based methods, such as RRT, RRG maintains cycles and improves connectivity by linking new samples to their nearest neighbors, thereby providing a denser approximation of the configuration space. This graph structure serves as the foundation for our method, and we briefly introduce two key functions that will be used in the following sections.
Steering Function
Given two states $s$ and $s^{\prime}$, the function $\text{Steer}(s,s^{\prime},\eta)$ returns a new state $s_{\text{new}}$ such that $s_{\text{new}}$ lies on the path connecting $s$ to $s^{\prime}$ and the distance $\|s-s_{\text{new}}\|_{2}\leq\eta$. The step size $\eta$ is a user-defined parameter that controls how far the path extends in a single step.
Sampling and connection
The sampling function generates a new state $s_{\text{new}}$ in the free space, and the connection function attempts to connect $s_{\text{new}}$ to its nearest neighbors $\mathcal{N}(s_{\text{new}})$ using a geometric trajectory. For RRG, if a connection is successful, the edge is added to the graph, forming a cycle when possible.
Given the graph $G=(V,E)$ constructed by RRG, if a node $s\in V$ lies within the goal region $\mathcal{S}_{G}$, we compute the optimal state trajectory $\mathbf{x}^{*}$ as a sequence of geometric states:
$\mathbf{x}^{*}=\{s_{0},s_{1},\ldots,s_{d}\},$
where $s_{0}$ is the start state, $s_{d}\in\mathcal{S}_{G}$, and $s_{i}\in V$ for $i=1,\ldots,d$. In most planning pipelines, Dijkstra’s algorithm [34] is employed to extract the shortest waypoint sequence from the constructed graph.
Beyond sampling-based graphs, grid-based graphs constitute another practical instantiation of our framework. In such cases, connectivity is predefined, and shortest paths are typically computed using algorithms such as A*. Additional details are provided in Appendix -A.
IV-B Motion Planning Guided Reward
After employing a graph-based motion planning method, we obtained a set of waypoints that span the entire graph. Based on this, we use an optimal path algorithm to compute the distance from each node to the goal region. We denote the optimal trajectory by $\mathbf{x^{*}}$, and the optimal is defined according to the length of trajectories. $s|\mathbf{x^{*}}$ denotes a state of trajectory $\mathbf{x^{*}}$. We define the distance function $Dist:\mathcal{S}\times\mathcal{S}\times\mathcal{X}\rightarrow[0,\infty]$ to represent the Euclidean distance between two states along a given trajectory, and define the cost function $Cost:\mathcal{S}\times\mathcal{X}\rightarrow[0,\infty]$ to represent the distance between one state and the destination $s_{d}$. Hence, the distance between two states along the optimal trajectory can be expressed as $Dist(s_{1},s_{2}|\mathbf{x^{*}})=|Cost(s_{1}|\mathbf{x^{*}})-Cost(s_{2}|\mathbf{x^{*}})|$. Consequently, the graph $G$ is now extended to $G=(V,E,Cost)$, where we slightly abuse the notation, denoting $Cost$ as distance between the current state and the goal.
The distance along the trajectory enables us to accurately quantify the actual path length, making it more applicable in cluttered environments compared to the Euclidean distance. Based on this distance, we design the motion-planning reward scheme to guide the RL agent. Given the difficulty of reaching an exact state in continuous space, we define the norm r-ball for loosen the criteria for determining whether the agent has reached specific states. Formally, it is defined by $Ball_{r}(s)=\left\{s^{\prime}\in\mathcal{S}\mid\|s-s^{\prime}\|_{2}\leq r\right\}$, where $r$ is the radius. The selection of the r-ball radius can influence the performance of the algorithm. For simplicity, we use the default $r=\frac{\eta}{2}$ based on the steering function in the following sections. We then define a progression function $D:\mathcal{S}\rightarrow[0,\infty]$ to determine whether the agent had moved toward the goal region and got closer to it. Formally,
$D(s)=\begin{cases}Cost\{s_{i}|\mathbf{x^{*}}\},&\text{if }x\in\text{Ball}_{r}(s_{i}\mid\mathbf{x^{*}})\\ \infty,&\text{otherwise}\end{cases}$ (2)
We adopt the definition of history from[35] with a slight modification $h_{t}=\{s_{1},a_{1},\dots,s_{t-1},a_{t-1}\}$ instead of $\{s_{1},a_{1},\dots,s_{t-1},a_{t-1},s_{t}\}$ , thereby the $D_{min}$ over a state-action sequence $\tau_{t}=h_{t}\bigcup\{s_{t},a_{t}\}$ is defined as $D_{min}(\mathbf{\tau})=\min_{s\in\mathbf{s}}{D(s)}$, where $\mathbf{s}$ denotes the state sequence in trajectory $\tau_{t}$. Ideally, we would check $D_{min}$ at each step to determine whether the agent has moved closer to the goal region, i.e., check if $D_{min}(\tau_{t})<D_{min}(\tau_{t-1})$, and assign a positive reward $R_{+}$ accordingly. However, this design introduces a non-Markovian reward issue. In other words, $r(h_{t},s_{t},a_{t})\neq r(h^{\prime}_{t},s_{t},a_{t})$ could happen with different history $h_{t}$ and $h^{\prime}_{t}$. To address this problem, we denote $\mathcal{S}^{\times}=\{(s,D_{min})\mid s\in\mathcal{S},\,D_{min}\in\mathbb{R}\}$, so that $D_{min}$ can be directly retrieved from the augmented state. Correspondingly, the augmented reward function is defined as $\tilde{r}:\mathcal{S}^{\times}\times\mathcal{A}\times\mathcal{S}^{\times}\to\mathbb{R}$,
$\tilde{r}_{t}(s_{t}^{\times},a_{t},s_{t+1}^{\times})=\left\{\begin{array}[]{ll}R_{-},&\text{if }s_{t+1}\text{ is in the obstacles},\\ R_{++},&\text{if }s_{t+1}\text{ is in the goal area},\\ R_{+},&\text{if }D_{min}^{t+1}<D_{min}^{t},\\ 0,&\text{otherwise.}\end{array}\right.$ (3)
where $R_{++}\gg R_{+}$ prioritizes goal achievement as the primary objective, while $R_{+}$ incentivizes exploratory progress toward the goal. We next analyze the properties of the reward function in this augmented state space.
Theorem 1.
(Markovian Analysis) By extending the state space to $\mathcal{S}^{\times}$, we ensure that the augmented reward function $\tilde{r}$ shown in Eq. (3) satisfies the Markov property. Formally,
$P\bigl{(}\tilde{r}_{t}\mid s_{0}^{\times},a_{0},\dots,s_{t}^{\times},a_{t},s_{t+1}^{\times}\bigr{)}\;=\;P\bigl{(}\tilde{r}_{t}\mid s_{t}^{\times},a_{t},s_{t+1}^{\times}\bigr{)}.$
for all time steps $t$, where $s_{t}^{\times}\in\mathcal{S}^{\times}$ denotes the augmented state including $D_{min}$.
This result ensures that our planning-augmented reward remains compatible with standard RL algorithms. In the remainder of this paper, we operate in the augmented state space $\mathcal{S}^{\times}$ to maintain this property. The detailed proof is presented in Appendix -B.
Proposition 1.
(Reward Performance) Optimizing the reward defined in Eq. (3) accelerates convergence and yields consistently high success rates in cluttered environments.
This proposition guarantees that the auxiliary rewards accelerate convergence without altering the original task objective. The proof is presented in Appendix -C. The learning objective aims to find the optimal policy by optimizing the parameters $\theta$:
$\theta^{*}=\arg\max_{\theta}J(\theta)=\mathbb{E}_{\pi_{\theta}}\left[\sum_{t=0}^{\infty}\gamma^{t}\cdot\tilde{r}_{t}(s_{t}^{\times},a_{t},s_{t+1}^{\times})\right]$
which corresponds to minimizing the following loss function:
$\theta^{*}=\arg\min_{\theta}\mathbb{E}_{(s^{\times},a,r,s^{\prime\times})\sim\mathcal{D}}\left[\left(Q_{\omega}(s,a)-y\right)^{2}\right],$
where $y$ is the target value, $\mathcal{D}$ is the replay buffer and $Q_{w}$ denotes the action-value function parameterized by $\omega$. It is worth noting that almost all reinforcement learning algorithms designed for cluttered environments adopt the actor-critic architecture [36, 17, 37, 18]. As noted in [36], actor-critic methods are sensitive to the quality and distribution of samples stored in the replay buffer, as effective optimization of the loss function depends heavily on informative and diverse experiences. Our proposed motion-planning based reward exhibits high spatial density, which facilitates exploration even under stochastic or suboptimal policies. Consequently, this promotes more stable and sample-efficient learning throughout training.
IV-C Infeasible Paths Analysis
Recalling Eq. (1), our method specifically addresses the challenge of unexpected disturbances in the simulation, which often render paths infeasible. In contrast to tree-based methods, which focus on planning a single path without accounting for its robustness or feasibility (e.g., with strong winds, as illustrated in Figure 4 (right) or other disturbances along the path), our approach leverages a graph-based motion planning strategy. By incorporating a comprehensive waypoint coverage across the entire map, the agent ensures that multiple potential paths are explored and evaluated during training.
We highlight that the robustness of our approach arises from the redundancy inherent in graph-based planning, which effectively mitigates the risk of path failure under disturbances. Instead of relying on a single predetermined trajectory, the agent benefits from multiple alternative paths towards the goal. As the density of waypoints and connectivity of the graph increase, the number of these alternative paths grows accordingly, greatly enhancing the likelihood that at least one safe and feasible path to the goal exists. This multiplicity ensures that even if certain paths become infeasible due to external perturbations, such as environmental disturbances, other viable paths remain available.
Example 2.
As shown in Figure 1 (right), compared to the tree-based method, our graph-based method constructs a dense waypoint network that enables comprehensive exploration and facilitates the learning of goal-reaching policies, even in cluttered environments and with unknown dynamics.
V Generalizing Initialization
With the graph-based waypoint covering the whole environment, we could naturally extend the training objective from single start point to arbitrary start points. This brings a desirable property, significant saving computational resource. In an autonomous system with tasks described by temporal logic formulas, if there are $N$ goal regions, in the worst case, if a neural network controller can only move from a specific starting point to a specific endpoint, the system would require $O(N^{2})$ distinct controllers. This indicates that in complex systems, the demand for controllers will increase dramatically as the number of destinations grows. However, as shown in figure 2 under our approach, the required number of controllers is only $O(N)$, i.e., only one controller is needed for each goal region. In this section, we will discuss how our method enables goal-reaching from arbitrary positions within the environment.
Figure 2: Generalize Initialization: We implement our method in the 3D maze environment and present both the top-down view (left) and the result after motion planning (right). The green point denotes the goal, the red point denotes the starting point, and the paths in different colors correspond to the optimal paths originating from different starting points.
V-A Discrete Region Partitioning
In RL training, it is often challenging to guarantee that a trained RL policy capable of completing a task in one initial region can achieve the same performance in another starting region [7, 38]. This limitation typically arises due to the incomplete exploration of the state space by the critic model, leading to underestimation issues. Here, inspired by [39] and [40], we first discretize the geometric space into a set of cell representations (as shown in figure 3).
Figure 3: Discrete Region Partitioning.
Let $\mathcal{E}\subset\mathbb{R}^{d}$ denote a continuous $d$-dimensional environment. We discretize $\mathcal{E}$ by partitioning it into a finite set of non-overlapping cells, each represented as a hypercube of fixed side length $\delta>0$. Formally, the discretized space can be expressed as a grid:
$\mathcal{E}_{\text{dis}}=\{c_{i_{1},i_{2},\dots,i_{d}}\mid i_{k}\in\mathbb{Z},\forall k=1,\dots,d\}$ (4)
where each cell $c_{i_{1},i_{2},\dots,i_{d}}$ is defined as
$c_{i_{1},i_{2},\dots,i_{d}}=\left\{x\in\mathbb{R}^{d}\mid i_{k}\delta\leq x_{k}<(i_{k}+1)\delta,\forall k=1,\dots,d\right\}.$ (5)
Here, $\delta$ is referred to as the spatial granularity, which determines the resolution of the discretization. The choice of $\delta$ has a certain impact on our method. However, for simplicity, we set $\delta$ as one-tenth of the smallest dimension length of the environment, providing a reasonable discretization resolution. Also we slightly abuse the notation $c_{i}$ as $c_{i_{1},i_{2},\dots,i_{d}}$, where $i$ is a shorthand for the tuple $(i_{1},i_{2},...,i_{d})$.
One intuitive advantage of partitioning the environment into discrete cells is that it significantly reduce the number of waypoints we need to iterate in the training process. Even in low-dimensional space with the same waypoint density, graph-based method generates significantly more waypoints and connections than tree-based methods. And this situation would be more extreme while we scale up the environment. By assigning each waypoint to its corresponding cell $c_{i}$, we can restrict the search within a much smaller region, significantly reducing the search space while maintaining efficient path planning. Therefore, here we extend the $G=(V,E,Cost)$ to $G=(V,E,Cost,Cell)$, where $Cell:\mathcal{S}\rightarrow\mathcal{E}_{dis}$ is a function mapping from geometry space to the discrete cell index.
V-B Weighted Sampled Starter
A common application of our approach is providing controllers for autonomous systems. Our method has already established reward guidance across every region of the environment. However, relying solely on randomly sampled starting points often fails to effectively explore all areas of the map and does not guarantee reachability from those regions to the goal. This limitation significantly impacts practical applications. For instance, in [41], verification results show that the set of controllers may fail to accomplish the specified task.
Recall the discretized state space be defined as $\mathcal{E}_{\text{dis}}=\{c_{i}\mid i\in\mathcal{I}\}$, where each $c_{i}$ represents a fixed-size grid cell. We define $N_{i}$ as the number of times the agent visits cell $c_{i}$ and compute the exploration coverage as $P_{i}=\frac{N_{i}}{\sum_{j\in\mathcal{I}}N_{j}+\epsilon}$ where $\epsilon$ is a small positive constant to prevent division by zero. To prioritize under-explored regions, we define the sampling weight: $w_{i}=\frac{1}{P_{i}+\epsilon}$, which is then normalized as $\tilde{w}_{i}=\frac{w_{i}}{\sum_{j\in\mathcal{I}}w_{j}}$. Using the normalized weights $\tilde{w}_{i}$, we sample a starting grid cell $c_{\text{start}}$ from a categorical distribution $c_{\text{start}}\sim\text{Categorical}(\mathcal{I},\tilde{w})$. Within the selected grid cell $c_{\text{start}}$, the exact starting state is uniformly sampled: $s_{\text{start}}\sim\mathcal{U}(c_{\text{start}})$.
As training progresses, visited states contribute to updating the visit counts of their respective grid cells, thereby gradually reducing the weight assigned to previously explored areas. Specifically, at the end of each training episode. For the trajectory $\tau=\{s_{0},\dots,s_{T},\}$ in the same episode, we first mapping from geometry space to cell space $\tau_{c}={Cell(s_{t}),\ \forall t\in[0,T]}$. Then we construct an unduplicated sets $\mathcal{C}_{visited}=\bigcup^{T}_{t=0}\{Cell(s_{t})\}$. The visit count for each explored cell $c_{i}$ is updated as: $N_{i}\leftarrow N_{i}+1,\quad\forall i\text{ where }c_{i}\in\mathcal{C}_{visited}$ This approach dynamically adjusts the start state distribution, ensuring thorough exploration and improving the coverage of the learned policy across diverse initializations.
VI Experimental Results
We validated our framework with different dynamic models. The experiments highlight that our method demonstrates superior performance in environments with disturbances where the environment changes after motion planning. We evaluated both our method and the baseline approaches using widely adopted reinforcement learning algorithms, i.e., PPO. We use high quality implementation from [42], mostly with default hyperparameters without further tuning. The results show that our approach is able to avoid potentially high-risk regions and significantly improves the success rate.
Figure 4: 2d and 3d navigation in environments with unmodeled disturbance, where light blue area, brown area and gray area represents wetted area, muddy area and strong wind area.
Figure 5: Comparison with other reward functions and validation of weighted sample start in 2D dynamics.
Figure 6: Comparison with other reward and the validation of weighted Sample Start in 3d dynamics
Baseline Approaches: The experiments were setup in both 2d and 3d cluttered environments where we compared our methods with others in different scales. We compare our method against the RRT-guided reward [24] and the binary reward to evaluate its ability to navigate cluttered environments with disturbances. Additionally, we examine the impact of the weighted sampled start strategy by comparing it with the standard approach without weighted sampling.
Quadrotor Model: We first implement the car-like model of Pybullet[43] physical engine shown in Figure 5. In this environment, we designed the obstacle layout to be highly cluttered, resembling a maze, while leaving a relatively simple path available-i.e., bypassing the maze from the upper corner.
We then blocked this simple path to demonstrate two key aspects of our method: it not only ensures successful goal-reaching but also naturally selects routes with lower potential risk. Compared to tree-based methods, which are sensitive to disturbances and unable to adjust their trajectory based on risk-being constrained to follow the pre-planned path-our approach exhibits significantly greater robustness. In this experiments, when we increased the disturbance level to the point that the RRT path was directly broken and the method failed entirely, our approach still performed reliably and successfully reached the goal. Meanwhile, we observed that the weighted sample start significantly improves the convergence speed compared to random initialization.
Autonomous Vehicle: We then implement the car-like model of Pybullet[43] physical engine shown in Figure 6. We randomly generated obstacles and introduced regions with strong winds, mud, and water along the path, representing lateral thrust, significant speed reduction, and impassable terrain, respectively. Similarly to the quadrotor scenario, when both binary reward and tree-based reward methods failed to complete the task, our framework still performed reliably. However, we did not observe a significant advantage of the weighted sample start over random start.
Success Certificate: We statistically run 200 trials of the learned policies for different selected start position and record the average success rates of both models, i.e., autonomous vehicle and quadrotor. The results are shown in Table I. We observe that in cluttered environments, assuming that all of the planned path is feasible and well tracked, the success rates of the binary reward is 0, while our method and tree-based reward achieve 100%; In cluttered environments which contain infeasible region, the success rates of all other baselines are 0, and our method achieves success rates near 100%. As a result, the robustness of the performance has improved significantly in response to environmental challenges.
TABLE I: Analysis of success rates.
Path Dynamic model Baseline rate Success rate
Feasible Quadrotor RRG 100%
Only RRT 100%
Binary 0%
Feasible Quadrotor RRG 100%
Only RRT 100%
Binary 0%
Contains Vehicle RRG 100%
Infeasible RRT 0%
Region Binary 0%
Contains Vehicle RRG 100%
Infeasible RRT 0%
Region Binary 0%
VII Conclusion
DRL suffers from inefficient exploration, especially in cluttered environments. Compared to random exploration, motion-planning-based methods can provide effective guidance for learning. However, tree-based planners that yield a single path may fail to adapt when environmental changes, limiting their guidance. This paper proposes a novel graph-based reward scheme that maintains a model-free setup while offering global guidance coverage to encourage optimal path exploration, and significantly reduce the number of controllers required. Our approach fully leverages the advantages of model-free RL and guarantees the goal-reaching performance for the original task. Future works include Sim2Real generalization and the extension to multi-agent collaboration.